OpenStudy (anonymous):

Solve the following system of equations for c: c + 6d = 31/2 d = 9 – 2c

OpenStudy (anonymous):

multiply 1st equation by 2 to get rid of the fractions 2c+12d=31 plug in d 2c+12(9-2c)=31 2c+81-24c=31 22c=50 c=50/22 =25/11

OpenStudy (anonymous):

Ohhh, thank you c:

OpenStudy (anonymous):

Wait. Wouldn't 12 times 9 be 108, not 81?

OpenStudy (anonymous):

yes you are right, I can't seem to multiply.

OpenStudy (anonymous):

Hah, it's okay. Thanks though!